OpenStudy (bsmith021):

Hi! I've honestly been trying to work out this problem for quite a while so far but I'm honestly really stuck on it ugh. Could someone possibly help me out! Thanks! The height of a triangle is four more than twice the length of the base. If the area of the triangle is 80 in.squared, find the length of the base and the height. (HINT: Formula for the area of a triangle is A=1/2 bh )

OpenStudy (august899):

Alright,So we know that the area is 80 in and the height is 2x+4 because it is 4 more than twice the length.

OpenStudy (bmk614):

To find the area of a triangle you do \[base*width*\frac{ 1 }{ 2 }\] Just set that equal to 80 and solve for x.

OpenStudy (vespasious):

Short answer: b=8, h=10 So... Hight is 2 times base plus 4. h=2b+4 If 1/2b*h = 80, that means b*h = 160. So if we know h=2b+4, and b*h = 160, we can determine... b*(2b+4) = 160. 2b^2 + 4b = 160 b^2 + 2b = 80 b^2 = 80-2b Notice that 8 squared = 64 2*8 = 16, 64+18 = 80, so 8 works for this answer. That means b=8. From there we can determine that height is 10, because b*h=160. b=8, h=10 Good luck!